# Large Language Models for Information Retrieval: A Survey

### v3

@@ 3.3.1 Prompting
TABLE I: Partial Examples of different prompting methods in query rewriting.
@@
$\bullet$ Zero-shot prompting. Zero-shot prompting involves instructing the model to generate texts on a specific topic without any prior exposure to training examples in that domain or topic. The model relies on its pre-existing knowledge and language understanding to generate coherent and contextually relevant expanded terms for original queries. Experiments show that zero-shot prompting is a simple yet effective method for query rewriting [109, 110, 99, 108, 98, 102].
@@
$\bullet$ Few-shot prompting. Few-shot prompting, also known as in-context learning, involves providing the model with a limited set of examples or demonstrations related to the desired task or domain [109, 110, 102, 86]. These examples serve as a form of explicit instruction, allowing the model to adapt its language generation to the specific task or domain at hand. Query2Doc [86] prompts LLMs to write a document that answers the query with some demo query-document pairs provided by the ranking dataset, such as MSMARCO [111] and NQ [112]. This work experiments with a single prompt. To further study the impact of different prompt designing, recent works [102] have explored eight different prompts, such as prompting LLMs to generate query expansion terms instead of entire pseudo documents and CoT prompting. There are some illustrative prompts in Table I. This work conducts more experiments than Query2Doc, but the results show that the proposed prompt is less effective than Query2Doc.
@@
$\bullet$ Chain-of-thought prompting. CoT prompting [45] is a strategy that involves iterative prompting, where the model is provided with a sequence of instructions or partial outputs [109, 102]. In conversational search, the process of query re-writing is multi-turn, which means queries should be refined step-by-step with the interaction between search engines and users. This process is naturally coincided with CoT process. As shown in 4, users can conduct the CoT process through adding some instructions during each turn, such as “Based on all previous turns, xxx”. While in ad-hoc search, there is only one-round in query re-writing, so CoT could only be accomplished in a simple and coarse way. For example, as shown in Table I, researchers add “Give the rationale before answering” in the instructions to prompt LLMs think deeply [102].
3.3.2 Fine-tuning
Fine-tuning is an effective approach for adapting LLMs to specific domains. This process usually starts with a pre-trained language model, like GPT-3, which is then further trained on a dataset tailored to the target domain. This domain-specific training enables the LLM to learn unique patterns, terminology, and context relevant to the domain, which is able to improve its capacity to produce high-quality query rewrites.
BEQUE [113] leverages LLMs for rewriting queries in e-commerce product searches. It designs three Supervised Fine-Tuning (SFT) tasks: quality classification of e-commerce query rewrites, product title prediction, and CoT query rewriting. To our knowledge, it is the first model to directly fine-tune LLMs, including ChatGLM [114, 68], ChatGLM2.0 [114, 68], Baichuan [115], and Qwen [116], specifically for the query rewriting task. After the SFT stage, BEQUE uses an offline system to gather feedback on the rewrites and further aligns the rewriters with e-commerce search objectives through an object alignment stage. Online A/B testing demonstrates the effectiveness of the method.
3.3.3 Knowledge Distillation
Although LLM-based methods have demonstrated significant improvements in query rewriting tasks, their practical implementation for online deployment is hindered by the substantial latency caused by the computational requirements of LLMs. To address this challenge, knowledge distillation has emerged as a prominent technique in the industry. In the QUILL [97] framework, a two-stage distillation method is proposed. This approach entails utilizing a retrieval-augmented LLM as the professor model, a vanilla LLM as the teacher model, and a lightweight BERT model as the student model. The professor model is trained on two extensive datasets, namely Orcas-I [117] and EComm [97], which are specifically curated for query intent understanding. Subsequently, a two-stage distillation process is employed to transfer knowledge from the professor model to the teacher model, followed by knowledge transfer from the teacher model to the student model. Empirical findings demonstrate that this knowledge distillation methodology surpasses the simple scaling up of model size from base to XXL, resulting in even more substantial improvements. In a recently proposed “rewrite-retrieve-read” framework [100], an LLM is first used to rewrite the queries by prompting, followed by a retrieval-augmented reading process. To improve framework effectiveness, a trainable rewriter, implemented as a small language model, is incorporated to further adapt search queries to align with both the frozen retriever and the LLM reader’s requirements. The rewriter’s refinement involves a two-step training process. Initially, supervised warm-up training is conducted using pseudo data. Then, the retrieve-then-read pipeline is described as a reinforcement learning scenario, with the rewriter’s training acting as a policy model to maximize pipeline performance rewards.
TABLE II: Summary of existing LLM-enhanced query rewriting methods. “Docs” and “KD” stand for document corpus and knowledge distillation, respectively.
Methods Target Data Generation
HyDE [97] Ad-hoc LLMs Prompting
Jagerman et al. [102] Ad-hoc LLMs Prompting
Query2Doc [86] Ad-hoc LLMs Prompting
Ma et al. [100] Ad-hoc LLMs Finetuning
PromptCase [103] Ad-hoc LLMs Prompting
GRF+PRF [99] Ad-hoc LLMs + Docs Prompting
GRM [96] Ad-hoc LLMs + Docs Prompting
InteR [98] Ad-hoc LLMs + Docs Prompting
LameR [108] Ad-hoc LLMs + Docs Prompting
CAR [107] Ad-hoc LLMs + Docs Prompting
QUILL [97] Ad-hoc LLMs + Docs KD & Finetuning
LLMCS [95] Conversational LLMs Prompting
CONVERSER [105] Conversational LLMs Prompting
Ye et al. [104] Conversational LLMs Prompting
@@ 3.4 Limitations
While LLMs offer promising capabilities for query rewriting, they also meet several challenges. Here, we outline two main limitations of LLM-based query rewriters.
@@ 3.4.1 Concept Drifts
When using LLMs for query rewriting, they may introduce unrelated information, known as concept drift, due to their extensive knowledge base and tendency to produce detailed and redundant content. While this can enrich the query, it also risks generating irrelevant or off-target results.
This phenomenon has been reported in several studies [118, 107, 113] These studies highlight the need for a balanced approach in LLM-based query rewriting, ensuring that the essence and focus of the original query are maintained while leveraging the LLM’s ability to enhance and clarify the query. This balance is crucial for effective search and IR applications.
@@ 3.4.2 Correlation between Retrieval Performance and Expansion Effects
Recently, a comprehensive study [119] conduct experiments on various expansion techniques and downstream ranking models, which reveals a notable negative correlation between retriever performance and the benefits of expansion. Specifically, while expansion tends to enhance the scores of weaker models, it generally hurts stronger models. This observation suggests a strategic approach: employ expansions with weaker models or in scenarios where the target dataset substantially differs in format from the training corpus. In other cases, it is advisable to avoid expansions to maintain clarity of the relevance signal.
@@ 4 Retriever
Unfortunately, there are some long-term challenges that hinder the advancement of retrieval models. First, user queries are usually short and ambiguous, making it difficult to precisely understand the user’s search intents for retrievers. Second, documents typically contain lengthy content and substantial noise, posing challenges in encoding long documents and extracting relevant information for retrieval models. Additionally, the collection of human-annotated relevance labels is time-consuming and costly. It restricts the retrievers’ knowledge boundaries and their ability to generalize across different application domains. Moreover, existing model architectures, primarily built on BERT [59], exhibit inherent limitations, thereby constraining the performance potential of retrievers. Recently, LLMs have exhibited extraordinary abilities in language understanding, text generation, and reasoning. This has motivated researchers to use these abilities to tackle the aforementioned challenges and aid in developing superior retrieval models. Roughly, these studies can be categorized into two groups, i.e., (1) leveraging LLMs to generate search data, and (2) employing LLMs to enhance model architecture.
@@ 4.1.1 Search Data Refinement
Typically, input queries consist of short sentences or keyword-based phrases that may be ambiguous and contain multiple possible user intents. Accurately determining the specific user intent is essential in such cases. Moreover, documents usually contain redundant or noisy information, which poses a challenge for retrievers to extract relevance signals between queries and documents. Leveraging the strong text understanding and generation capabilities of LLMs offers a promising solution to these challenges. As yet, research efforts in this domain primarily concentrate on employing LLMs as query rewriters, aiming to refine input queries for more precise expressions of the user’s search intent. Section 3 has provided a comprehensive overview of these studies, so this section refrains from further elaboration. In addition to query rewriting, an intriguing avenue for exploration involves using LLMs to enhance the effectiveness of retrieval by refining lengthy documents. This intriguing area remains open for further investigation and advancement.
Figure 5: Two typical frameworks for LLM-based data augmentation in the retrieval task (right), along with their prompt examples (left). Note that the methods of relevance label generation do not treat questions as inputs but regard their generation probabilities conditioned on the retrieved passages as soft relevance labels.
@@ 4.1.2 Training Data Augmentation
How to apply LLMs for data augmentation? In the scenario of IR, it is easy to collect numerous documents. However, the challenging and costly task lies in gathering real user queries and labeling the relevant documents accordingly. Considering the strong text generation capability of LLMs, many researchers [120, 122] suggest using LLM-driven processes to create pseudo queries or relevance labels based on existing collections. These approaches facilitate the construction of relevant query-document pairs, enlarging the training data for retrieval models. According to the type of generated data, there are two mainstream approaches that complement the LLM-based data augmentation for retrieval models, i.e., pseudo query generation and relevance label generation. Their frameworks are visualized in Figure 5. Next, we will give an overview of the related studies.
$\bullet$ Pseudo query generation. Given the abundance of documents, a straightforward idea is to use LLMs for generating their corresponding pseudo queries. One such illustration is presented by inPairs [120], which leverages the in-context learning capability of GPT-3. This method employs a collection of query-document pairs as demonstrations. These pairs are combined with a document and presented as input to GPT-3, which subsequently generates possible relevant queries for the given document. By combining the same demonstration with various documents, it is easy to create a vast pool of synthetic training samples and support the fine-tuning of retrievers on specific target domains. Recent studies [121] have also leveraged open-sourced LLMs, such as Alpaca-LLaMA and tk-Instruct, to produce sufficient pseudo queries and applied curriculum learning to pre-train dense retrievers. To enhance the reliability of these synthetic samples, a fine-tuned model (e.g., a monoT5-3B model fine-tuned on MSMARCO [122]) is employed to filter the generated queries. Only the top pairs with the highest estimated relevance scores are kept for training. This “generating-then-filtering” paradigm can be conducted iteratively in a round-trip filtering manner, i.e., by first fine-tuning a retriever on the generated samples and then filtering the generated samples using this retriever. Repeating these EM-like steps until convergence can produce high-quality training sets [123]. Furthermore, by adjusting the prompt given to LLMs, they can generate queries of different types. This capability allows for a more accurate simulation of real queries with various patterns [124].
@@
4.2 Employing LLMs to Enhance Model Architecture
Leveraging the excellent text encoding and decoding capabilities of LLMs, it is feasible to understand queries and documents with greater precision compared to earlier smaller-sized models [59]. Researchers have endeavored to utilize LLMs as the foundation for constructing advanced retrieval models. These methods can be grouped into two categories, i.e., dense retrievers and generative retrievers.
@@ 4.2.1 Dense Retriever
In addition to the quantity and quality of the data, the representative capability of models also greatly influences the efficacy of retrievers. Inspired by the LLM’s excellent capability to encode and comprehend natural language, some researchers [130, 131, 132] leverage LLMs as retrieval encoders and investigate the impact of model scale on retriever performance.
General Retriever. Since the effectiveness of retrievers primarily relies on the capability of text embedding, the evolution of text embedding models often has a significant impact on the progress of retriever development. In the era of LLMs, a pioneer work is made by OpenAI [130]. They view the adjacent text segments as positive pairs to facilitate the unsupervised pre-training of a set of text embedding models, denoted as cpt-text, whose parameter values vary from 300M to 175B. Experiments conducted on the MS MARCO [111] and BEIR [128] datasets indicate that larger model scales have the potential to yield improved performance in unsupervised learning and transfer learning for text search tasks. Nevertheless, pre-training LLMs from scratch is prohibitively expensive for most researchers. To overcome this limitation, some studies [131, 133] use pre-trained LLMs to initialize the bi-encoder of dense retriever. Specifically, GTR [133] adopts T5-family models, including T5-base, Large, XL, and XXL, to initialize and fine-tune dense retrievers. RepLLaMA [131] further fine-tunes the LLaMA model on multiple stages of IR, including retrieval and reranking. For the dense retrieval task, RepLLaMA appends an end-of-sequence token “<</s>>” to the input sequences, i.e., queries or documents, and regards its output embeddings as the representation of queries or documents. The experiments confirm again that larger model sizes can lead to better performance, particularly in zero-shot settings. Notably, the researchers of RepLLaMA [131] also study the effectiveness of applying LLaMA in the reranking stage, which will be introduced in Section 5.1.3.
Task-aware Retriever. While the aforementioned studies primarily focus on using LLMs as text embedding models for downstream retrieval tasks, retrieval performance can be greatly enhanced when task-specific instructions are integrated. For example, TART [132] devises a task-aware retrieval model that introduces a task-specific instruction before the question. This instruction includes descriptions of the task’s intent, domain, and desired retrieved unit. For instance, given that the task is question-answering, an effective prompt might be “Retrieve a Wikipedia text that answers this question. {question}”. Here, “Wikipedia” (domain) indicates the expected source of retrieved documents, “text” (unit) suggests the type of content to retrieve, and “answers this question” (intent) demonstrates the intended relationship between the retrieved texts and the question. This approach can take advantage of the powerful language modeling capability and extensive knowledge of LLMs to precisely capture the user’s search intents across various retrieval tasks. Considering the efficiency of retrievers, it first fine-tunes a TART-full model with cross-encoder architecture, which is initialized from LLMs (e.g., T0-3B, Flan-T5). Then, a TART-dull model initialized from Contriever [134] is learned by distillating knowledge from the TART-full.

### v4

@@ 3.3.1 Prompting
TABLE II: Examples of different prompting methods in query rewriter.
@@
$\bullet$ Zero-shot prompting. Zero-shot prompting involves instructing the model to generate texts on a specific topic without any prior exposure to training examples in that domain or topic. The model relies on its pre-existing knowledge and language understanding to generate coherent and contextually relevant expanded terms for original queries. Experiments show that zero-shot prompting is a simple yet effective method for query rewriter [78, 98, 82, 85, 84, 76].
@@
$\bullet$ Few-shot prompting. Few-shot prompting, also known as in-context learning, involves providing the model with a limited set of examples or demonstrations related to the desired task or domain [78, 98, 76, 65]. These examples serve as a form of explicit instruction, allowing the model to adapt its language generation to specific tasks or domains. Query2Doc [65] prompts LLMs to write a document that answers the query with some demo query-document pairs provided by the ranking dataset, such as MSMARCO [99] and NQ [100]. This work experiments with a single prompt. To further study the impact of different prompt designing, recent works [76] have explored eight different prompts, such as prompting LLMs to generate query expansion terms instead of entire pseudo documents and CoT prompting. Some illustrative prompts are shown in Table II. The experiments validate that Query2Doc is more effective than many other prompt-based methods.
@@
$\bullet$ Chain-of-thought prompting. CoT prompting [97] is a strategy that involves iterative prompting, where the model is provided with a sequence of instructions or partial outputs [78, 76]. In conversational search, the process of query rewriting is multi-turn, which means queries should be refined step-by-step with the interaction between search engines and users. This process naturally coincides with the CoT process. As shown in Figure 3, users can conduct the CoT process by adding some instructions during each turn, such as “Based on all previous turns, xxx”. While in ad-hoc search, there is only one round in query rewriting, so CoT could only be accomplished in a simple and coarse way. For example, as shown in Table II, researchers add “Give the rationale before answering” in the instructions to prompt LLMs to think deeply [76].
3.3.2 Supervised Fine-tuning
Prompting methods directly leverage LLMs’ strong capabilities to expand or rewrite queries. Though prompting method is effective, LLMs are not naturally designed for query rewriting task. To further tailor LLMs for this task, supervised fine-tuning (SFT) has emerged as a promising approach. A crucial aspect of this methodology is the creation of an appropriate training dataset. The process of gathering this dataset varies significantly depending on the application scenario.
In the context of e-commerce search, a wealth of supervised training data for query rewriting is naturally available. This data, sourced from the previous-generation rewriting policies of the e-commerce system, significantly simplifies the construction of the SFT dataset [81].
Conversely, in an ad-hoc retrieval scenario, the acquisition of query rewrite training data is often a challenge. To address this issue, researchers usually employ implicit feedback and reinforcement learning to train the query rewriter.
3.3.3 Reinforcement Learning
Query rewriters typically serve as intermediaries for retrieval systems, implying that they do not operate independently, and there is no specific loss attributed solely to the query rewriting process. In such a context, Reinforcement Learning (RL) presents an alternative training mechanism for the query rewriter. Query rewriters can receive feedback signals from ranking models [88] or downstream LLM readers [80]. For instance, ranking scores can be utilized to construct good-bad pairs for DPO [101] training. Ma et al. [80] proposed a method that first generates answers from LLMs and then uses the QA evaluation results as training signals. In a similar manner, BEQUE [81] introduced an offline feedback system capable of providing a quality score for each query based on the set of products it retrieved. Such an application of Reinforcement Learning mechanisms aligns the objective of query rewriters more closely with the goals of downstream tasks, thereby enhancing the overall performance of the system.
@@ 3.4 Limitations
Despite the potential of LLMs in query rewriting, they still suffer from several limitations. We discuss two primary challenges that arise with their use in this context.
@@ 3.4.1 Concept Drifts
One significant issue is the introduction of unrelated information, or concept drift, which may occur due to the LLM’s vast knowledge base and propensity to generate detailed yet sometimes redundant content. While this can potentially enrich the query, it may also lead to irrelevant or off-target results. This issue has been reported in various studies [94, 87, 81]. These works emphasize the necessity for a balanced approach to LLM-based query rewriting, maintaining the core and focus of the original query while utilizing the LLM’s capabilities to enhance and clarify it. This balance is critical for effective search and IR applications.
@@ 3.4.2 Correlation between Retrieval Performance and Expansion Effects
A recent comprehensive study [102] has investigated various expansion techniques and downstream ranking models, revealing a significant negative correlation between retrieval performance and expansion benefits. Specifically, expansion tends to improve the scores of weaker models but adversely affects stronger ones. This finding suggests a strategic approach: only using expansions with weaker models or when the target dataset significantly varies in format from the training corpus. In other situations, it may be more beneficial to refrain from expansions to preserve the clarity of the relevance signal.
@@ 4 Retriever
Unfortunately, there are some long-term challenges that hinder the advancement of retrieval models. First, user queries are usually short and ambiguous, making it difficult to precisely understand the user’s search intents for retrievers. Second, documents typically contain lengthy content and substantial noise, posing challenges in encoding long documents and extracting relevant information for retrieval models. Additionally, the collection of human-annotated relevance labels is time-consuming and costly. It restricts the retrievers’ knowledge boundaries and their ability to generalize across different application domains. Moreover, existing model architectures, primarily built on BERT [55], exhibit inherent limitations, thereby constraining the performance potential of retrievers. Recently, LLMs have exhibited extraordinary abilities in language understanding, text generation, and reasoning. This has motivated researchers to use these abilities to tackle the aforementioned challenges and aid in developing superior retrieval models. Roughly, these studies can be categorized into two groups: (1) leveraging LLMs to generate search data, and (2) employing LLMs to enhance model architecture.
@@ 4.1.1 Search Data Refinement
Typically, input queries consist of short sentences or keyword-based phrases that may be ambiguous and contain multiple possible user intents. Accurately determining the specific user intent is essential in such cases. Moreover, documents usually contain redundant or noisy information, which poses a challenge for retrievers to extract relevance signals between queries and documents. Leveraging the strong text understanding and generation capabilities of LLMs offers a promising solution to these challenges. As yet, research efforts in this domain primarily concentrate on employing LLMs as query rewriters, aiming to refine input queries for more precise expressions of the user’s search intent. Section 3 has provided a comprehensive overview of these studies, so this section refrains from further elaboration. In addition to query rewriter, an intriguing avenue for exploration involves using LLMs to enhance the effectiveness of retrieval by refining lengthy documents. This intriguing area remains open for further investigation and advancement.
Figure 4: Three typical frameworks for LLM-based data augmentation in the retrieval task (right), along with their prompt examples (left). Note that the methods of relevance label generation do not treat questions as inputs but regard their generation probabilities conditioned on the retrieved passages as soft relevance labels.
@@
Gecko [110] few-shot Unkown Relevant query - Fixed
@@ 4.1.2 Training Data Augmentation
How to apply LLMs for data augmentation? In the scenario of IR, it is easy to collect numerous documents. However, the challenging and costly task lies in gathering real user queries and labeling the relevant documents accordingly. Considering the strong text generation capability of LLMs, many researchers [103, 105] suggest using LLM-driven processes to create pseudo queries or relevance labels based on existing collections. These approaches facilitate the construction of relevant query-document pairs, enlarging the training data for retrieval models. According to the type of generated data, there are three mainstream approaches that complement the LLM-based data augmentation for retrieval models, i.e., pseudo query generation, relevance label generation, and complete example generation. Their frameworks are visualized in Figure 4.
$\bullet$ Pseudo query generation. Given the abundance of documents, a straightforward idea is to use LLMs for generating their corresponding pseudo queries. One such illustration is presented by inPairs [103], which leverages the in-context learning capability of GPT-3. This method employs a collection of query-document pairs as demonstrations. These pairs are combined with a document and presented as input to GPT-3, which subsequently generates possible relevant queries for the given document. By combining the same demonstration with various documents, it is easy to create a vast pool of synthetic training samples and support the fine-tuning of retrievers on specific target domains. To enhance the diversity of generated examples, Gecko [110] prompts LLMs to first generate a task description, and then generate pseudo queries according to the task. Considering the false negative problems, it further develops an LLM-based positive and negative mining strategy to discover potential relevant and hard negative documents from the corpus for generated queries, which significantly enhance the retrieval performance. Recent studies [104] have also leveraged open-sourced LLMs, such as Alpaca-LLaMA and tk-Instruct, to produce sufficient pseudo queries and applied curriculum learning to pre-train dense retrievers. To enhance the reliability of these synthetic samples, a fine-tuned model (e.g., a monoT5-3B model fine-tuned on MSMARCO [105]) is employed to filter the generated queries. Only the top pairs with the highest estimated relevance scores are kept for training. This “generating-then-filtering” paradigm can be conducted iteratively in a round-trip filtering manner, i.e., by first fine-tuning a retriever on the generated samples and then filtering the generated samples using this retriever. Repeating these EM-like steps until convergence can produce high-quality training sets [106]. Furthermore, by adjusting the prompt given to LLMs, they can generate queries of different types. This capability allows for a more accurate simulation of real queries with various patterns [107].
@@
$\bullet$ Complete example generation. Recent study [112] further investigates approaches to directly utilize LLMs to generate synthetic queries and documents, hence providing tremendous diverse training examples across varied tasks and languages. This work proposes a two-stage generation pipeline, where the first stage prompts the LLM to brainstorm various retrieval tasks, and then the second stage generates corresponding “(query, positive document, negative document)” triplets to build synthetic training data. Researchers further control the length, languages, and semantic relationships of queries and documents to produce diverse training samples. The generated triplets, after post-processing, e.g., de-duplication and JSON-format filtering, are used as training examples for optimizing dense retrievers.
@@
4.2 Leveraging LLMs as Retrievers’ Backbone
Thanks to the superior text representation capability, LLMs excel at comprehending the underlying semantics of queries and documents. Therefore, it becomes increasingly popular to apply such large-scale models as the backbone of text retrievers, leading to substantial improvements over the conventional methods based on smaller-sized models [55].
@@ 4.2.1 Dense Retriever
The application of LLMs brings about two major impacts on dense retrieval. On one hand, it advances the ongoing progress of the existing methods, making substantial improvements in terms of both in-domain accuracy and out-of-domain generalizability. On the other hand, it extends the boundaries of current methods, introducing new capabilities such as instruction following and in-context learning.
Improved Existing Capacities. Dense retrieval needs to fine-tune pre-trained text encoders such that queries and documents can be transformed into semantic-rich embeddings. Therefore, the downstream retrieval performance can benefit from the utilization of stronger foundations. With preliminary progresses achieved by early forms of pre-trained models (e.g., BERT [55] and T5 [58]), people moved on to take advantage of LLMs for dense retrieval. In this direction, a pioneer work is made by OpenAI, where a series of GPT models were fine-tuned towards text and code representation [115]. It is the first time where decoding-only transformers were effectively applied for dense retrieval. Importantly, it empirically validates that the retrieval performance can be consistently improved from the increased model size and embedding dimension. Besides, the LLM-based retrievers also exhibit superior generalizability [116, 117], as notable improvements can be achieved for not only the targeted scenario but also a variety of general tasks beyond the fine-tuned domain.
Recently, the development of LLM-based dense retrievers have gotten dramatically promoted as powerful LLMs, e.g., LLaMA [35], Vicuna [118], Mistral [119], Phi [120], and Gemma [121], are made publicly available. Remarkably, RepLLaMA [122], the first fine-tuned embedder on top of open-source LLM (LLaMA-2-7B), brings forth major improvements on a variety of benchmarks, including MSMARCO passage/doc retrieval [99] and BEIR [113]. Despite extra computation costs due to the expanded model scale, the first-stage retrieval accuracy with RepLLaMA alone already surpasses the multi-stage retrieval accuracy resulted from conventional methods, indicating the its potential value for real-world application. After that, people make exploration of other alternatives for dense retrieval, where additional improvements are continually achieved with adoption of more advanced LLMs [112, 123, 124]. To date, LLM-based embedders have dominated all major text retrieval benchmarks, e.g., currently, the leading methods on MTEB [125] are all back-ended by LLMs.
In addition to the above methods which directly fine-tune LLMs, there are also parallel works on adapting generic LLMs as better foundations for dense retrieval. For example, Llama2Vec [126] performs post pre-training of LLaMA-2 with two new pretext tasks: EBAE (embedding based auto-encoding) and EBAR (embedding based auto-regression). With moderate scale of training on unlabeled corpus, it results in substantial improvements of retrieval performance over the basic Llama-2 model. Besides, NV-Embed modifies LLM’s architecture by introducing latent attention layer and bidirectional attention [127]. Both modifications contribute to the improved performance on MTEB benchmark. Despite the above preliminary progresses, there are still many open challenges about LLM-based embedders, such as efficiency and adaptability, which need to be addressed in the future.
Introducing New Capacities. Compared to conventional methods that use small-scale pre-trained models, LLM-based embedders introduce new capabilities that enhance the usability and accuracy of dense retrieval. One notable example is their ability to follow instructions, allowing LLM-based embedders to be trained for various semantic matching tasks based on user demands. For instance, an LLM-based embedder can perform document retrieval when prompted with “retrieve relevant docs for the input question”, and can be adapted for duplicate question retrieval with the prompt “retrieve questions with the same meaning as input”. Although BERT-based retrievers are also fine-tuned to follow instructions [128, 129], they do not support unseen instructions as effectively as LLM-based embedders [130]. ChatRetriever [131] further leverages dialog-based instruction tuning to build LLM-based conversational embedders, enhancing their conversational retrieval capabilities. In addition to instructions, the LLM-based embedders can also be adapted through in-context learning, where the retrieval function can be updated by demonstration examples of user’s interested tasks [132]. Another advantage of LLM-based embedders is their length-generalizable capacity, which allows them to effectively handle much longer texts than those in their training examples [130]. This makes it possible to manage retrieval applications across various text lengths while maintaining a feasible training cost.